OpenStudy (anonymous):

someone wanna help me? x squared = 12 ...solve using square root property?

OpenStudy (anonymous):

Take square root of both sides.

OpenStudy (anonymous):

so its x=3 rad 2?

OpenStudy (anonymous):

What is the square root over the two or the three?

OpenStudy (anonymous):

idk?

OpenStudy (anonymous):

That's not a good answer.

OpenStudy (anonymous):

idk how to solve this problem? i dont know what the sqauare root of 12 is. all i no is that it is 3 rad 2......what else do i need to do?

OpenStudy (anonymous):

You keep the 12 under the radical and break it down: 2.2.3 all under the square root sign.

OpenStudy (anonymous):

well...so its not 3 square root of 2?

OpenStudy (anonymous):

Right. Actually it is sq rt 4=2

OpenStudy (anonymous):

how is it the square root of 4?

OpenStudy (anonymous):

sq rt of 12 is equal to sq rt of (4.3)

OpenStudy (anonymous):

Not to confuse you, I was just correcting your statement\[2\sqrt{3}\]

OpenStudy (anonymous):

so is square root of (x-5) ^ 2 = sqaure root of 2 does x =6?

OpenStudy (anonymous):

What?! You lost me.

OpenStudy (anonymous):

ok i feel dumb. ok solve this using the square root property x-5 ^ 2 =2

OpenStudy (anonymous):

i got 6

OpenStudy (anonymous):

Same process: sq rt of both sides.

OpenStudy (anonymous):

ok i did and got x-5 = 1

OpenStudy (anonymous):

sq rt of 2 is not 1

OpenStudy (anonymous):

o...well 2/2 = 1

OpenStudy (anonymous):

That is correct, but not part of this problem.

OpenStudy (anonymous):

ugh ok so what is the square root of 2?

OpenStudy (anonymous):

Sq rt of 2 is sq rt of 2.

OpenStudy (anonymous):

but i have to simplyfy ? whats the answer of the equartion then? just x-5 = sqr 2?

OpenStudy (anonymous):

You can get a number by using calculator, but since you are studying sq rt you can leave it in that form, same way you left sq rt of 3 in previous problem

OpenStudy (anonymous):

do you no how to compelte the square?

OpenStudy (anonymous):

Yes, it's that thingy where you divide by 2, then square it

OpenStudy (anonymous):

ok x ^ 2 - 8x = 3

OpenStudy (anonymous):

slve by completing the squarue

OpenStudy (anonymous):

You take that thing in front of x, divide it by 2, square it, add it to each side.

OpenStudy (anonymous):

??? can u write it out?

OpenStudy (anonymous):

theres nothing infront of x

OpenStudy (anonymous):

Looks like -8 to me.

OpenStudy (anonymous):

oo yeah i got u

OpenStudy (anonymous):

so it would be 16x ^ 2 - 4x = 19?

OpenStudy (anonymous):

You did more than my instructions said, the x^2 -8x remains part of the equation, the thing you calculated is added but they are apples and oranges you just write it in.

OpenStudy (anonymous):

so how would you write it?

OpenStudy (anonymous):

\[x ^{2}-8x+16=3+16\]

OpenStudy (anonymous):

ok a couple more.... how would you do 1+5i/3-2i

OpenStudy (anonymous):

Have to look up my old notes, been a while.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

You have to do conjugation: multiply by (3+2i)/(3+2i)

OpenStudy (anonymous):

wait how can u multiply and divide at the same time??

OpenStudy (anonymous):

its 2AM...im sorry im sooo tired but this is due at 8AM

OpenStudy (anonymous):

You have to manipulate your original problem in order to have something to work with. The choice of conjugation is an off-shoot of the bottom term of your problem. In math you can't change things but you can do tricks, for example if you multiply by 1, you have not changed anything. The thing I told you to multiply is equal to 1.

OpenStudy (anonymous):

ok so how do u solve the problem?

OpenStudy (anonymous):

You use the FOIL method multiply top The top looks like this (1+5i)(3+2i) The bottom is (3-2i)(3+2i)

OpenStudy (anonymous):

i got 13 + 17i

OpenStudy (anonymous):

I assume you are working on the top only, recall FOIL method: First, Outside, Inside, Lasts. Be careful to add only like terms, add apples to apples and oranges to oranges.

OpenStudy (anonymous):

that what i did cuz u said the bottom = 1

OpenStudy (anonymous):

Well, the bottom is not equal to 1, the whole term I introduced is equal to 1. You skip to many steps. Take baby steps, produce the terms of FOIL, eg First is1x3=3

OpenStudy (anonymous):

ok .... this is the answer i got..of its wrong just tell me the answer....... 13+17i / 13-12i

OpenStudy (anonymous):

You can't get that, my first term is 3

OpenStudy (anonymous):

3+ 2i +15i +10i^2 which i^2 = -1 over 9-6i -6 - 4i^2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

That's it you got rid of some of the i's but there is still one left in top, must check notes again.

OpenStudy (anonymous):

ok how do you do 6/cube root of 29?

OpenStudy (anonymous):

Not much you can do with that, if you are work with cube roots like you were with square roots.

OpenStudy (anonymous):

o snap i wrote that wrong... i meant 6/cube root of 2a ^ 2 sorry!

OpenStudy (anonymous):

I am looking over my notes for complex numbers, you are doing everything right; seems the point is to at least get it out of the bottom. Break it down as much you can, final answer would have i in it.

OpenStudy (anonymous):

i need that rationalized

OpenStudy (anonymous):

For these, they just want you to move square root out of the the bottom. So the same process as complex number, you multiply by\[\sqrt[3]{2a ^{2}}/\sqrt[3]{2a ^{2}}\]

OpenStudy (anonymous):

so it would be 6 cube root 2a ^2 over ????

OpenStudy (anonymous):

If you multiply cube root of something by cube root of something you would get something. So the answer is whatever is under radical.

OpenStudy (anonymous):

2a? is the top right?

OpenStudy (anonymous):

2a^2, yes top is right.

OpenStudy (anonymous):

ok last problem!! square root of 2x+11 =5

OpenStudy (anonymous):

You do the opposite of earlier problems: square both sides.

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

You skipped to many steps. Jumping in shower.

OpenStudy (anonymous):

im going to bed night